# The multivalued inverse of a quasiregular map is a quasiregular curve

### v1

@@ 4.2 Pull-back by multi-valued maps
The proof is identical to Lemma 6.1, Lemma 6.2 and Proposition 6.3 in [pan-onn21]. The only place a minor argument is needed is at the beginning of Lemma 6.1, where it is stated that $\omega=\mathrm{d}\tau$ for a suitable (n1)(n-1)-form $\tau$. We need to ensure that $\tau$ can be chosen $S_{d}$-invariant. But this follows from Lemma 4.2: $\omega=P_{S_{d}}\omega=\mathrm{d}(P_{S_{d}}\tau)$, and $\tau^{\prime}=P_{S_{d}}\tau$ is $S_{d}$-invariant. The rest of the proof(s) proceed exactly as in [pan-onn21]. Note that $\mathrm{d}(f^{*}\tau^{\prime})=f^{*}\omega$ (a fact needed in the proof in [pan-onn21]) holds by Theorem 1.1. ∎
5 Multivalued inverse of QR map as a QC map
@@
5.1 Image of multivalued inverse
@@
A lower bound for $\frac{|U_{\accentset{\leftarrow}{f}\circ f}(x,r)|}{r^{n}}$, together with the LLC-property, would imply a local $(1,1)$-Poincaré inequality for $\Omega_{f}$ by a deep result of Semmes [sem96].
@@
In particular, $\Omega_{f}$ is an $n$-dimensional differentiability space, cf. [teri-toni-enrico, Theorem 1.1].
In the proof we use the terminology and techniques from [teri-syl24]. Since these are used only in this proof we do not recall them here but refer the reader instead to [teri-syl24].
@@
for path families in $\Omega$, cf. Proposition 5.6 and [iko-luc-pas24, Lemma 3.6].
@@
We claim that $(B_{0},F_{0}^{-1})$ is both a Cheeger chart and an $n$-weak chart. Since $F_{0}$ is metrically differentiable everywhere and has $(N)$ and $(N^{-1})$, it is not difficult to see that $(B_{0},F_{0})$ is a Cheeger chart. [^1] Indeed, a direct calculation yields that $(B_{0},F_{0}^{-1})$ is a weak Cheeger chart. Since $B_{0}$ is open, it follows that $(B_{0},F_{0}^{-1})$ is a Cheeger chart. See [iko-luc-pas24] for the terminology. It also follows that, if $(B_{0},F_{0})$ is $n$-independent (in the sense of [teri-syl24]) then it is maximal, and thus an $n$-weak chart.
@@
Now since $(B_{0},F_{0}^{-1})$ is an $n$-dimensional Cheeger chart and $n$-weak chart, it follows from [teri-syl24, Theorem 1.11] and [teri-toni-enrico, Theorem 1.3] that $\operatorname{Lip}g=|Dg|_{n}$ $\mathcal{H}^{n}$-a.e. on $\Omega_{f}$ for any $g\in\operatorname{LIP}(\Omega_{f})$. This completes the proof. ∎
5.2 Quasiconformality of multi-valued inverse
@@
For a Sobolev map $f\in N^{1,n}(Z,W)$ with Lusin’s properties $(N)$ and $(N^{-1})$ between $n$-rectifiable spaces satisfying the infinitesimal $n$-PI, the minimal upper gradient $|Df|$ of $f$ coincides $\mathcal{H}^{n}$-a.e. $x\in Z$ with $\|D_{x}f\|_{op}$, where $D_{x}f$ is the approximate differential of $f$ at $x$, and the volume Jacobian of [wil12b] coincides with $\mathbf{J}f=\mathbf{J}(D_{x}f)$. Thus the inequality $|Df|^{n}\leq K\ \mathbf{J}f$ is equivalent with the modulus inequality
@@
Now, by rectifiability and the infinitesimal Poincaré inequality we have that Gigli’s co-tangent module $L^{n}(T^{*}Z)$ coincides with the $n$-integrable sections $\Gamma_{n}(T^{*}Z)$ of the co-tangent bundle given by the rectifiable atlas, cf. [teri-toni-enrico, Theorems 1.2 and 1.3], $Z=X,Y$. The claim now follows from [iko-luc-pas24, Theorem 3.13]. Indeed, Using the above identification (and the fact that $h$ has Lusin’s properties $(N)$ and $(N^{-1})$ we have that the pull-back operator $h^{*}$ constructed in [iko-luc-pas24, Theorem 3.4] coincides with $(Dh)^{*}:T^{*}Y\to T^{*}X$, where $Dh:TX\to TY$ is the bundle map given by the a.e. defined approximate differential of $h$. Since
@@
$(Dh)^{*}$ is an isomorphism, proving condition (i) in [iko-luc-pas24, Theorem 3.13], while condition (ii) in [iko-luc-pas24, Theorem 3.13] is satisfied since $h^{-1}\in N_{loc}^{1,p}(Y,X)$. ∎
@@
We will prove the claim by induction on the degree of $f$. Denote by $I(k)$ the induction hypothesis: $\accentset{\leftarrow}{f}^{-1}\in N^{1,n}_{loc}(U_{f},\mathbb{R}^{n})$ and $|D\accentset{\leftarrow}{f}^{-1}|\leq\frac{C(f)}{H\circ\accentset{\leftarrow}{f}^{-1}}$ for every quasiregular map $f:U\to\mathbb{R}^{n}$ of degree $\leq k$ from an open set $U\subset\mathbb{R}^{n}$, where $C(f)=(K_{I}(f)K_{O}(f))^{1/n}$. When $k=1$, the claim is classical. Indeed, a locally homeomorphic quasiregular map is quasiconformal [ric93, Chapter III, Corollary 3.8], and its inverse is quasiconformal with the correct bound [ric93, Chapter II, Corollary 6.5]. Thus $I(1)$ holds.
@@
Note that $G_{\varepsilon}=f\circ b$ in a neighbourhood of $A$, $G_{\varepsilon}=\accentset{\leftarrow}{f}^{-1}$ in $\Omega_{f}\setminus B(A,\varepsilon)$, and $G_{\varepsilon}\in N_{loc}^{1,n}(\Omega_{f},\mathbb{R}^{n})$. [^2] Indeed, for all points in $B(A,\varepsilon)\setminus A$, $G_{\varepsilon}$ is a composition of maps in $N^{1,n}$.
@@
We close the paper with the proof of Theorem 1.4.

### v2

@@ 5. Multivalued inverse of QR map as a QR curve
The proof is identical to Lemma 6.1, Lemma 6.2 and Proposition 6.3 in [21]. The only place a minor argument is needed is at the beginning of Lemma 6.1, where it is stated that $\omega=\mathrm{d}\tau$ for a suitable (n1)(n-1)-form $\tau$. We need to ensure that $\tau$ can be chosen $S_{d}$-invariant. But this follows from Lemma 4.2: $\omega=P_{S_{d}}\omega=\mathrm{d}(P_{S_{d}}\tau)$, and $\tau^{\prime}=P_{S_{d}}\tau$ is $S_{d}$-invariant. The rest of the proof(s) proceed exactly as in [21]. Note that $\mathrm{d}(f^{*}\tau^{\prime})=f^{*}\omega$ (a fact needed in the proof in [21]) holds by Theorem 1.1. ∎
The following theorem states that the multi-valued inverse of a QR map is a quasiminimizer of the $n$-energy and the parametrized volume, respectively, which are defined for $u\in W^{1,n}_{loc}(\Omega,\mathcal{A}_{d}(\mathbb{R}^{n}))$ and $V\subset\Omega$ as follows.
$E_{n}(u;V):=\int_{V}\|Du\|^{n}\mathrm{d}x,\quad A(u;V):=\int_{V}\mathbf{J}u\ \mathrm{d}x.$
Theorem 5.4.
Let $\Omega\subset\mathbb{R}^{n}$ be a domain and $f:\Omega\to\mathbb{R}^{n}$ a proper QR map. If $u\in W_{loc}^{1,n}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n}))$ is such that $u=\accentset{\leftarrow}{f}$ a.e. outside a compact set $K\subset\Omega$, then
$E_{n}(\accentset{\leftarrow}{f};K)$ $\leq d^{n/2-1}K_{I}E_{n}(u;K)$
$A(\accentset{\leftarrow}{f};K)$ $\leq K_{I}A(u;K)$
Recall that $K_{I}=K_{I}(f)$ is the inner distortion of $f$.
Proof.
Let $\tau$ be an (n1)(n-1)-form on $\mathcal{A}_{d}(\mathbb{R}^{n})$ such that $\omega=\mathrm{d}\tau$. Since $u=\accentset{\leftarrow}{f}$ a.e. on $\Omega\setminus K$, we have that $u^{*}\omega=\accentset{\leftarrow}{f}^{*}\omega$ and $u^{*}\tau=\accentset{\leftarrow}{f}^{*}\tau$ a.e. on $\Omega\setminus K$.
Set $K_{\varepsilon}:=\{x:\ \operatorname{dist}(K,x)\leq\varepsilon\}$, and note that $K_{\varepsilon}$ is a finite perimeter set for a.e. $\varepsilon$. Moreover $u^{*}\tau=\accentset{\leftarrow}{f}^{*}\tau$ a.e. on $\partial^{*}K_{\varepsilon}$ for a.e. $\varepsilon$ by a Fubini-type argument. Thus, by a generalized Stoke’s theorem, we have
$\int_{K_{\varepsilon}}u^{*}\omega=\int_{K_{\varepsilon}}\mathrm{d}(u^{*}\tau)=\int_{\partial^{*}K_{\varepsilon}}u^{*}\tau=\int_{\partial^{*}K_{\varepsilon}}\accentset{\leftarrow}{f}^{*}\tau=\int_{K_{\varepsilon}}\mathrm{d}(\accentset{\leftarrow}{f}^{*}\tau)=\int_{K_{\varepsilon}}\accentset{\leftarrow}{f}^{*}\omega$ (5.4)
for a.e. $\varepsilon>0$. (Note that we use Theorem 1.1 in the first and last equalities above.) Indeed, $\llbracket K_{\varepsilon}\rrbracket$ is a normal $n$-current on $\mathbb{R}^{n}$ and $\partial\llbracket K_{\varepsilon}\rrbracket=\llbracket\partial^{*}K_{\varepsilon}\rrbracket$ whenever $K_{\varepsilon}$ has finite perimeter, and in this case we have by [15, Theorem 1.1] the following: for each $v\in W^{1,n}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n}))$ the identity $\partial(v_{\ast}\llbracket K_{\varepsilon}\rrbracket)=v_{\ast}(\partial\llbracket K_{\varepsilon}\rrbracket)=v_{\ast}(\llbracket\partial^{*}K_{\varepsilon}\rrbracket)$ holds a.e. $\varepsilon$. This implies
$\int_{K_{\varepsilon}}\mathrm{d}(v^{*}\beta)=\int_{\partial^{*}K_{\varepsilon}}v^{*}\beta$
for smooth (n1)(n-1)-forms $\beta$ on $\mathcal{A}_{d}(\mathbb{R}^{n})$ (cf. also [15, Lemma 5.3]), which in turn implies (5.4) for a.e. $\varepsilon$. Using (5.4) we may estimate
$E_{n}(\accentset{\leftarrow}{f};K)$ $\leq\int_{K_{\varepsilon}}\|D\accentset{\leftarrow}{f}\|^{n}\mathrm{d}x\leq d^{n/2-1}K_{I}\int_{K_{\varepsilon}}\star\accentset{\leftarrow}{f}^{*}\omega=d^{n/2-1}K_{I}\int_{K_{\varepsilon}}\accentset{\leftarrow}{f}^{*}\omega$
$=d^{n/2-1}K_{I}\int_{K_{\varepsilon}}u^{*}\omega\leq d^{n/2-1}K_{I}E_{n}(u;K_{\varepsilon})$
for a.e. $\varepsilon>0$. The last inequality above follows from the pointwise inequalities $\star u^{*}\omega\leq\mathbf{J}u\leq\|Du\|^{n}$, cf. Lemma 5.1. Since $\mathcal{L}^{n}(K_{\varepsilon}\setminus K)\to 0$ as $\varepsilon\to 0$, we obtain
$E_{n}(\accentset{\leftarrow}{f};K)\leq d^{n/2-1}K_{I}E_{n}(u;K).$
Similarly,
$A(\accentset{\leftarrow}{f};K)\leq\int_{K_{\varepsilon}}\mathbf{J}\accentset{\leftarrow}{f}\mathrm{d}x\leq K_{I}\int_{K_{\varepsilon}}\accentset{\leftarrow}{f}^{*}\omega=K_{I}\int_{K_{\varepsilon}}u^{*}\omega\leq K_{I}A(u;K_{\varepsilon})$
(cf. Remark 5.2 and Lemma 5.1) yielding
$A(\accentset{\leftarrow}{f};K)\leq K_{I}A(u;K)$
at the limit $\varepsilon\to 0$. This completes the proof. ∎
Remark 5.5.
The use of pull-back theory and the generalized Stokes theorem could also be used to give another proof of Theorem 5.3. Indeed, it yields a reverse Hölder inequality without a Caccioppoli inequality, and the higher integrability is then a standard consequence of a theorem of Gehring. We omit the details.
We end this section with a proof of Corollary 1.3, which is essentially a consequence of Theorems 5.3 and 5.4.
Proof of Corollary 1.3.
Claim (i) follows immediately from Theorem 5.3, Corollary A.2(i) and the observation that $g_{f}=b\circ\accentset{\leftarrow}{f}$.
To prove (ii), let $h\in W^{1,n}_{loc}(f(\Omega),\mathbb{R}^{n})$ be such that $\{h\neq g_{f}\}$ is essentially contained in a compact subset $K\subset f(\Omega)$. A straightforward calculation shows that if $R>0$, then the multivalued map $R\accentset{\leftarrow}{f}$, defined by $R\accentset{\leftarrow}{f}(x)=\sum_{y\in f^{-1}(x)}\iota(f,x)\llbracket Rx\rrbracket$, is an energy quasi-minimizer with the same constant as $\accentset{\leftarrow}{f}$ and $\|D(R\accentset{\leftarrow}{f})\|=R\|D\accentset{\leftarrow}{f}\|$. Now consider the multivalued function
$h_{R}=\mathcal{L}^{-1}\circ(R\cdot h,\pi_{\mathcal{Z}}\circ\accentset{\leftarrow}{f})$
Then $\{h_{R}\neq\accentset{\leftarrow}{f}\}$ is essentially contained in $K$, and by Theorem 5.4 and the discussion above we have
$\int_{K}R^{n}\|D\accentset{\leftarrow}{f}\|^{n}\mathrm{d}x\leq d^{n/2-1}K_{I}\int_{K}\|Dh_{R}\|^{n}\mathrm{d}x.$
By Corollary A.2(ii) this implies
$\int_{K}d^{n/2}R^{n}\|Dg_{f}\|^{n}\leq d^{n/2-1}K_{I}\int_{K}(dR^{2}\|Dh\|^{2}+\|D(\pi_{\mathcal{Z}}\circ\accentset{\leftarrow}{f})\|^{2})^{n/2}\mathrm{d}x.$
Dividing by $R^{n}$ and taking the limit $R\to+\infty$ we obtain
$\int_{K}\|Dg_{f}\|^{n}\mathrm{d}x\leq d^{n/2-1}K_{I}\int_{K}\|Dh\|^{n}\mathrm{d}x.$
6. Multivalued inverse of QR map as a QC map
@@
6.1. Image of multivalued inverse
@@
A lower bound for $\frac{|U_{\accentset{\leftarrow}{f}\circ f}(x,r)|}{r^{n}}$, together with the LLC-property, would imply a local $(1,1)$-Poincaré inequality for $\Omega_{f}$ by a deep result of Semmes [27].
@@
In particular, $\Omega_{f}$ is an $n$-dimensional differentiability space, cf. [14, Theorem 1.1]. In the proof we use the terminology and techniques from [3]. Since these are used only in this proof we do not recall them here but refer the reader instead to [3].
@@
for path families in $\Omega$, cf. Proposition 6.6 and [12, Lemma 3.6].
@@
We claim that $(B_{0},F_{0}^{-1})$ is both a Cheeger chart and an $n$-weak chart. Since $F_{0}$ is metrically differentiable everywhere and has $(N)$ and $(N^{-1})$, it is not difficult to see that $(B_{0},F_{0})$ is a Cheeger chart. [^2] Indeed, a direct calculation yields that $(B_{0},F_{0}^{-1})$ is a weak Cheeger chart. Since $B_{0}$ is open, it follows that $(B_{0},F_{0}^{-1})$ is a Cheeger chart. See [12] for the terminology. It also follows that, if $(B_{0},F_{0})$ is $n$-independent (in the sense of [3]) then it is maximal, and thus an $n$-weak chart.
@@
Now since $(B_{0},F_{0}^{-1})$ is an $n$-dimensional Cheeger chart and $n$-weak chart, it follows from [3, Theorem 1.11] and [14, Theorem 1.3] that $\operatorname{Lip}g=|Dg|_{n}$ $\mathcal{H}^{n}$-a.e. on $\Omega_{f}$ for any $g\in\operatorname{LIP}(\Omega_{f})$. This completes the proof. ∎
6.2. Quasiconformality of multi-valued inverse
@@
For a Sobolev map $f\in N^{1,n}(Z,W)$ with Lusin’s properties $(N)$ and $(N^{-1})$ between $n$-rectifiable spaces satisfying the infinitesimal $n$-PI, the minimal upper gradient $|Df|$ of $f$ coincides $\mathcal{H}^{n}$-a.e. $x\in Z$ with $\|D_{x}f\|_{op}$, where $D_{x}f$ is the approximate differential of $f$ at $x$, and the volume Jacobian of [30] coincides with $\mathbf{J}f=\mathbf{J}(D_{x}f)$. Thus the inequality $|Df|^{n}\leq K\ \mathbf{J}f$ is equivalent with the modulus inequality
@@
Now, by rectifiability and the infinitesimal Poincaré inequality we have that Gigli’s co-tangent module $L^{n}(T^{*}Z)$ coincides with the $n$-integrable sections $\Gamma_{n}(T^{*}Z)$ of the co-tangent bundle given by the rectifiable atlas, cf. [14, Theorems 1.2 and 1.3], $Z=X,Y$. The claim now follows from [12, Theorem 3.13]. Indeed, Using the above identification (and the fact that $h$ has Lusin’s properties $(N)$ and $(N^{-1})$ we have that the pull-back operator $h^{*}$ constructed in [12, Theorem 3.4] coincides with $(Dh)^{*}:T^{*}Y\to T^{*}X$, where $Dh:TX\to TY$ is the bundle map given by the a.e. defined approximate differential of $h$. Since
@@
$(Dh)^{*}$ is an isomorphism, proving condition (i) in [12, Theorem 3.13], while condition (ii) in [12, Theorem 3.13] is satisfied since $h^{-1}\in N_{loc}^{1,p}(Y,X)$. ∎
@@
We will prove the claim by induction on the degree of $f$. Denote by $I(k)$ the induction hypothesis: $\accentset{\leftarrow}{f}^{-1}\in N^{1,n}_{loc}(U_{f},\mathbb{R}^{n})$ and $|D\accentset{\leftarrow}{f}^{-1}|\leq\frac{C(f)}{H\circ\accentset{\leftarrow}{f}^{-1}}$ for every quasiregular map $f:U\to\mathbb{R}^{n}$ of degree $\leq k$ from an open set $U\subset\mathbb{R}^{n}$, where $C(f)=(K_{I}(f)K_{O}(f))^{1/n}$. When $k=1$, the claim is classical. Indeed, a locally homeomorphic quasiregular map is quasiconformal [26, Chapter III, Corollary 3.8], and its inverse is quasiconformal with the correct bound [26, Chapter II, Corollary 6.5]. Thus $I(1)$ holds.
@@
Note that $G_{\varepsilon}=f\circ b$ in a neighbourhood of $A$, $G_{\varepsilon}=\accentset{\leftarrow}{f}^{-1}$ in $\Omega_{f}\setminus B(A,\varepsilon)$, and $G_{\varepsilon}\in N_{loc}^{1,n}(\Omega_{f},\mathbb{R}^{n})$. [^3] Indeed, for all points in $B(A,\varepsilon)\setminus A$, $G_{\varepsilon}$ is a composition of maps in $W^{1,n}$.
@@
We close the section with the proof of Theorem 1.4.